1. A mass breaks suddenly into two parts, masses M and m, with speeds V and v respectively and carrying a total kinetic energy K. If M / m = R =, what fractions of K does each of the masses carry? (Think about significant figures.)

a) M carries _____ K.

b) m carries _____ K.

2. A machine gun fires bullets, each with mass m = 55 g at a speed of u = 450 m.s^-1. The gun fires 5.0 bullets per second. Let's use it as a rocket engine. What thrust does it produce?

3. Let's combine momentum and collisions with some of the other things we've learned. A small block of mass M = 0.201 kg hangs on the end of a light, in extensible string, length R = 25 cm. A small dart (of mass m=0.10 kg) traveling in the horizontal direction collides with and remains fixed in the block. What is the minimum speed v of the dart such that the combined object completes a circular path around the support point of the block? How fast must the block be traveling at the top?
1
Expert's answer
2020-03-10T11:24:46-0400

1) From the conservation of momentum:


"mv=MV=RmV\to v=RV"

"K=0.5mv^2+0.5Rm\frac{v^2}{R^2}=0.5mv^2\left(1+\frac{1}{R}\right)"

a)


"K_m=\frac{1}{\left(1+\frac{1}{R}\right)}K=\frac{R}{R+1}K"

b)


"K_M=\left(1-\frac{R}{R+1}\right)K=\frac{1}{R+1}K"

2)


"F=mu\frac{N}{t}"

"F=(0.055)(450)(5)=120\ N"

3) From the conservation of momentum:


"mv=(m+M)u"

At the top:


"m\frac{u_t^2}{R}=mg"

"u_t=\sqrt{gR}=\sqrt{(9.8)(0.25)}=1.57\frac{m}{s}"

From the conservation of energy:


"u^2=u_t^2+2Rg"

"u=\sqrt{3gR}"

So,


"mv=(m+M)\sqrt{3gR}"

"0.1v=(0.1+0.201)\sqrt{3(9.8)(0.25)}"

"v=8.2\frac{m}{s}"